OpenStudy (anonymous):

47 is 20% of what number

OpenStudy (anonymous):

Two ways: 1) 47 is 20% Divide both by 2 to get 23.5 is 10% Multiply both by 10 235 is 100% Answer: 235 OR 2) 47/20% = 47 / 20 * 100 = 2.35 * 100 = 235

OpenStudy (anonymous):

To check: 235 x 20% = 235 x 20 / 100 = 4700 / 100 = 47